Given a standard normal distribution with mean =62 and standard deviation = 44, find P(X>92)


1
Expert's answer
2022-01-26T14:46:26-0500
"P(X>92)=1-P(X\leq92)="

"=1-P(Z\leq\dfrac{92-\mu}{\sigma})=1-P(Z\leq\dfrac{92-62}{44})"

"\approx 1-P(Z\leq0.681818)\approx0.247677"